# FineMolTex: Towards Fine-grained Molecular Graph-Text Pre-training

### v2

@@ 3.2 Pre-training Tasks
Contrastive Alignment. For coarse-grained alignment at the molecule level, we align the graph-text pairs from the same molecules and contrast the pairs from different molecules, which can be achieved by optimizing the following loss.
@@
$+\left.\log\dfrac{\exp(\cos(\mathbf{z_{t_{0}}},\mathbf{z_{m_{0}}})/\tau)}{\exp(\cos(\mathbf{z_{t_{0}}},\mathbf{z_{m_{0}}})/\tau)+\sum_{m_{0}^{\prime}}\exp(\cos(\mathbf{z_{t_{0}}},\mathbf{z_{m_{0}^{\prime}}})/\tau)}\right],$ (9)
@@
Masked Multi-modal Modeling. For fine-grained alignment at the motif level, we mask some of the tokens and predict their labels. Based on the fragmented motifs of all molecules in the pre-training dataset, we construct a motif dictionary that includes motifs along with their labels and frequency. Then we randomly mask approximately 20% of the motif tokens that have neither too high nor too low a frequency in the motif dictionary, as well as 15% of the word tokens. The token embeddings of the motifs and words are updated utilizing $\ell_{\text{trm}_{M}}$ and $\ell_{\text{trm}_{T}}$ transformer layers, respectively. Subsequently, information from the two modalities is integrated via our cross-attention layer. This entire process is iterated for $\ell$ times.
@@
$L_{\text{pre}}=\beta\sum_{i}\text{CE}(\hat{y}_{m_{i}},{y}_{m_{i}})+\alpha\sum_{j}\text{CE}(\hat{y}_{t_{j}},{y}_{t_{j}}),$ (10)
@@
4 Experiments
In this section, we conduct extensive experiments to demonstrate the effectiveness of FineMolTex. Before evaluating, we first conduct the two pre-training tasks on the PubChemSTM dataset [26], which includes 281K graph-text pairs from PubChem [16]. Each molecular graph is paired with a textual description that elaborates on its chemical and physical properties or highlights its high-level bioactivities. Details of the pre-training data and process can be found in Appendix C.1.1 and  C.4.
@@ 4 Experiments
4.1 Generalization to Unseen Molecules (RQ1)
Table 1: Accuracy (%±$\pm\sigma$) of graph-text retrieval task on DrugBank-Pharmacodynamics.
@@ 4.1 Generalization to Unseen Molecules (RQ1)
FineMolTex 95.86±$\mathbf{\pm 0.34}$ 91.95±$\mathbf{\pm 0.06}$ 85.80±$\mathbf{\pm 0.05}$ 95.80±$\mathbf{\pm 0.06}$ 92.18±$\mathbf{\pm 0.12}$ 85.01±$\mathbf{\pm 0.32}$
Table 2: Accuracy (%±$\pm\sigma$) of graph-text retrieval task on molecule-ATC.
@@
FineMolTex 75.43±$\mathbf{\pm 0.15}$ 60.66±$\mathbf{\pm 0.08}$ 49.45±$\mathbf{\pm 0.24}$ 75.22±$\mathbf{\pm 0.12}$ 60.29±$\mathbf{\pm 0.04}$ 48.42±$\mathbf{\pm 0.15}$
@@
Datasets and Baselines. We utilize DrugBank-Pharmacodynamics, molecule-ATC, and DrugBank-Description [26] extracted from the DrugBank database [44] for evaluation. These datasets include molecular graphs and their chemical descriptions. Details of the datasets can be found in Appendix  C.1.2. We compare with five multimodal molecular models: KV-PLM [53], MolCA [27], MoMu-S [35], MoMu-K [35], and MoleculeSTM [26]. Specifically, KV-PLM uses SMILES to represent the structure of the molecule, while others use graph structures.
@@
Results. We report the results on the first two datasets in Tables LABEL:tab:pha and  LABEL:tab:atc, and defer those on DrugBank-Description to Appendix D.1 due to space limit. We make the following observations. 1) Across different values of $T$, FineMolTex consistently outperforms the baselines that neglect motif-level knowledge. The superior performance demonstrates that fine-grained motif-level knowledge facilitates generalization to unseen molecules, which likely contain seen motifs. 2) FineMolTex maintains strong performance in both directions (given graph, and given text). The symmetry further indicates that the embedding spaces of both modalities are well-aligned and similarly well-learned. 3) We observe that KV-PLM, which utilizes SMILES to capture molecular structures, is less effective than other models employing graphs, consistent with previous findings [26] that 2D graph structure is more expressive than 1D SMILES.
4.2 Application to Motif-Centered Tasks (RQ2)
Figure 3: Hit ratios of 12 text-based molecule editing tasks.
@@ 4.2 Application to Motif-Centered Tasks (RQ2)
Figure 4: Visual analysis of the output molecules of MoleculeSTM and Motif-MolTex on 4 text-based molecule editing tasks. Differences between the input and output molecule of FineMolTex are highlighted in red and green circles. Lower LogP indicates higher water solubility.
@@
To answer RQ2, we employ a zero-shot text-based molecule editing task, which is highly relevant to practical applications including catalyst design and targeted drug discovery. Specifically, we utilize FineMolTex to collaborate with a molecule generation module, following the design in [26], to modify a specified molecule according to a text prompt. Hence, motif-level knowledge is essential for this task, as the model needs to replace certain motifs with others that are related to specific properties and names as indicated in the text prompt. We defer the technical details to Appendix B. We randomly sample 200 molecules from ZINC [14], and select 12 text prompts, including 8 prompts pertaining to physical properties [26], and 4 based on the names of the motifs. We utilize MoleculeSTM, MoMu, and MolCA as baselines.
@@
Evaluation. We employ different methods to assess whether the generated molecules satisfy the two types of prompts. For the 8 prompts on physical properties, we employ three measures: LogP, QED, and tPSA, which measures solubility [21], drug-likeness [4], and permeability [8], respectively. We consider the editing to be successful if the difference in measurements between the input and output molecules exceeds a specified threshold $\Delta$, which we have set to 00 following one of the settings in literature [26]. For the 4 prompts based on motif names, we use RDKit [20] to verify the presence of the indicated motifs. For all 12 prompts, we report the hit ratio: the proportion of generated molecules that meet our expectations.
@@
Results. As shown in Figure 3, FineMolTex shows superior performance on these prompts, especially on the 4 prompts with motif names. Notably, we achieve a relative gain of up to 230% over the best-performing baseline, demonstrating that FineMolTex has an advanced understanding of motif-level knowledge. We also visualize the output molecules of MoleculeSTM and FineMolTex in Figure 4. It can be observed that while MoleculeSTM produces incorrect molecules, FineMolTex accurately generates the intended molecules. Specifically, when prompted to generate molecules that are soluble in water, FineMolTex successfully creates molecules with lower LogP than the input molecule, as lower LogP indicates higher water solubility. Similarly, when prompted to generate molecules with hydroxyl groups or chlorine atoms, FineMolTex correctly does so. These results confirm that FineMolTex possesses a deeper understanding of motif-level knowledge, thereby enhancing the generative capabilities. More visual results can be seen in Appendix D.2.
4.3 Application to Single-Modality Task (RQ3)
@@
Table 3: Downstream results (%±$\pm\sigma$) on eight binary classification datasets from MoleculeNet.
@@
MolCA 70.0$\pm$0.5 77.2$\pm$0.5 64.5$\pm$0.8 63.0$\pm$1.7 89.5$\pm$0.7 72.1±$\pm 1.3$ 77.2±$\pm 0.6$ 79.8$\pm$0.5 74.2
MoleculeSTM 70.0±$\pm 0.5$ 76.9±$\pm 0.5$ 65.1±$\pm 0.4$ 61.0±$\pm 1.05$ 92.5±$\pm 1.1$ 73.4±$\pm 2.9$ 77.0±$\pm 1.8$ 80.8±$\pm 1.3$ 74.6
FineMolTex 71.0±$\mathbf{\pm 0.4}$ 77.1±$\mathbf{\pm 0.5}$ 66.0±$\mathbf{\pm 1.2}$ 65.0±$\mathbf{\pm 2.3}$ 92.7±$\mathbf{\pm 0.8}$ 76.2±$\mathbf{\pm 1.2}$ 78.9±$\mathbf{\pm 0.6}$ 84.6±$\mathbf{\pm 1.4}$ 76.4
While FineMolTex can simultaneously utilize pre-trained knowledge from both graphs and texts, we also verify its effectiveness on single-modality tasks, namely, molecular property prediction tasks. We use MoleculeNet [46] as the dataset, which only provides molecular graphs as input without texts. More specifically, there are eight binary classification tasks, and we report ROC-AUC for evaluation. More detailed dataset descriptions are provided in Appendix C.1.3.
Baselines. We compare FineMolTex against nine baselines, including 1) five pre-trained GNN models: AttrMasking [12], ContextPred [12], InfoGraph [36], MolCLR [43], and GraphMVP [25]; 2) three graph-text multimodal models: MoMu-S [35], MoMu-K [35], and MoleculeSTM [26], and 3) one SMILES-text multimodal model: KV-PLM [53].
Results. As shown in Table  3, FineMolTex consistently outperforms all baselines, achieving relative gains of 3.2%, 2.4%, and 4.7% on SIDER, MUV, and BACE, respectively, compared to the best baseline. The promising performance of FineMolTex indicates that it implicitly utilizes pre-trained knowledge from the text modality even when the input consists solely of graphs. Additionally, KV-PLM exhibits a notable performance gap from other models, due to its use of 1D SMILES strings for molecular structure and a smaller pre-training dataset.
4.4 Analysis of Learned Fine-grained Knowledge (RQ4)
(a) Visualization of motif tokens and word tokens using $t$-SNE. Triangles denote word tokens; circles denote motif tokens.
(b) Explaination of the prediction of certain masked motifs based on text tokens utilizing LIME.
Figure 5: Case studies validating motif-level knowledge learned by our model.
Visualization of Motif and Word Embeddings. To evaluate whether FineMolTex captures fine-grained alignment knowledge, we select motif and word tokens from 1,500 graph-text pairs in the PubChemSTM dataset, excluding meaningless words such as “this” and “a”. In total, we visualize 3,469 motif tokens and 6,914 text tokens with $t$-SNE [29] in Figure 5(a), where triangles denote text tokens, and circles denote motif tokens, with different colors indicating various labels. To examine the details of the tokens, we zoom into several regions in the figure, retaining only the colors and legends of the tokens we are interested in. For brevity, we utilize SMILES to represent the motif structures. We observe that text and motif tokens corresponding to each other are also close in the embedding space. For instance, in the orange frame, the word “sulf” is close to the motif tokens “OS”, “CCSSCC”, and “CC1=CSC(C)=N1”, all of which represent sulfides. In the pink frame, the word “ammonium” is close to the motif tokens “[NH2+]=O”, “C=[NH2+]”, and “[NH3+]O”, which are related to “ammonium.” In the blue frame, the word “poison” is adjacent to the motifs “[AsH3]” and “O[AsH2]”, which are poisonous. These results demonstrate that FineMolTex learns the connections between motifs and their chemical names and properties, thereby significantly enhancing its expressiveness.
Predictions Based on Fine-grained knowledge. To further verify that FineMolTex can utilize the learned fine-grained alignment knowledge for predictions, we utilize Local Interpretable Model-Agnostic Explanation (LIME) [31], a well-established tool that can explain the predictions of certain masked motifs based on text tokens. By perturbing the input text, LIME observes how the model’s predictions change with variations in the input text. Then, LIME fits these perturbed texts and the prediction results to an interpretable model such as a linear model. This approach allows us to quantify the significance of each text token in predicting the motifs, thereby revealing the fine-grained knowledge learned by FineMolTex. The results are shown in Figure 5(b) (with more cases in Appendix D.4), where text tokens with higher interpretive weights are more crucial for predictions, and thus more relevant to the masked motifs. Specifically, the word with the highest interpretive weights in (1) is “triazoles”, which directly is to the name of the masked motif. In (2), the word “aniline” refers to another motif that is very similar to the masked motif, and “aromatic” is the property related to the masked motif. In (3), the masked motif is a kind of “organonitrogen”. These findings demonstrate that FineMolTex has effectively acquired motif-level knowledge.
4.5 Ablation Study for Masking and Cross-Attention Layers (RQ5)
To thoroughly explore the impact of the key components in FineMolTex, we compare to several variants, including w/o mmm, which drops the masked multimodal modeling task altogether; motif mask only, which only mask motif tokens; word mask only, which only mask word tokens; w/o cross-attention, which excludes cross-attention layers. We evaluate these variants on the graph-text retrieval task used in RQ1, on two datasets with $T=4$. As reported in Table LABEL:tab:abla, FineMolTex consistently surpasses the other variants. Specifically, without the masked multimodal modeling task, “w/o mmm” fails to capture fine-grained knowledge at all, resulting in the poorest performance. “motif mask only” and “word mask only” outperform “w/o mmm,” because they still enable some level of fine-grained knowledge learning by either predicting motifs based on word tokens or predicting word tokens based on motif tokens. However, they are less effective than FineMolTex, which masks both words and motifs for mutual alignment. Lastly, without the cross-attention layers, “w/o cross-attention” cannot integrate token embeddings from different modalities, thereby hampering its ability to effectively learn fine-grained knowledge. These observations demonstrate the effectiveness of each component.
Table 4: Ablation study (%±$\pm\sigma$) on molecule-ATC and DrugBank-Pharmacodynamics.
@@ 4.5 Ablation Study for Masking and Cross-Attention Layers (RQ5)
w/o mmm 70.33±$\pm 0.28$ 70.86±$\pm 0.16$ 92.56±$\pm 0.10$ 92.12±$\pm 0.25$
motif mask only 72.85±$\pm 0.46$ 71.15±$\pm 0.24$ 93.45±$\pm 0.44$ 92.98±$\pm 0.18$
word mask only 74.27±$\pm 0.07$ 72.11±$\pm 0.13$ 94.86±$\pm 0.32$ 93.88±$\pm 0.15$
w/o cross-attention 71.44±$\pm 0.32$ 70.92±$\pm 0.18$ 92.66±$\pm 0.08$ 92.85 ±$\pm 0.24$
FineMolTex 75.43±$\mathbf{\pm 0.15}$ 75.22±$\mathbf{\pm 0.12}$ 95.86±$\mathbf{\pm 0.34}$ 95.80±$\mathbf{\pm 0.06}$
5 Conclusions, Limitations and Broader Impact

### v5

@@ 3.2. Pre-training Tasks
Contrastive Alignment. For coarse-grained alignment at the molecule level, we align the graph-text pairs from the same molecules and contrast the pairs from different molecules, which can be achieved by optimizing the following loss:
@@
(9) $+\left.\log\dfrac{\exp(\cos(\mathbf{z_{t_{0}}},\mathbf{z_{m_{0}}})/\tau)}{\exp(\cos(\mathbf{z_{t_{0}}},\mathbf{z_{m_{0}}})/\tau)+\sum_{m_{0}^{\prime}}\exp(\cos(\mathbf{z_{t_{0}}},\mathbf{z_{m_{0}^{\prime}}})/\tau)}\right],$
@@
Masked Multi-modal Modeling. For fine-grained alignment at the motif level, we selectively mask the important tokens and predict their labels. The token embeddings of the motifs and words are updated using $\ell_{\text{trm}_{M}}$ and $\ell_{\text{trm}_{T}}$ transformer layers, respectively. Subsequently, information from the two modalities is integrated via our cross-attention layer. This entire process is iterated for $\ell$ times.
@@
(10) $L_{\text{pre}}=\beta\sum_{i}\text{CE}(\hat{y}_{m_{i}},{y}_{m_{i}})+\alpha\sum_{j}\text{CE}(\hat{y}_{t_{j}},{y}_{t_{j}}),$
@@
3.3. Importance-based Masking
There is a large proportion of noisy tokens that lack fine-grained knowledge. For example, motif tokens such as “C”, and word tokens like “an” and “this” fail to provide meaningful information. If they are used for capturing fine-grained alignment knowledge, they may negatively impact the training process.
To address this issue, we propose the importance score to identify and mask tokens that contain fine-grained alignment knowledge. Specifically, as shown in Figure  2, we observe that important motif tokens, such as $m_{5}$, tend to contribute more to the global molecule token $z_{m_{0}}$, meaning they have a larger weight during aggregation. The same applies to the global text token $z_{t_{0}}$. Intuitively, the contrastive alignment task encourages the global token to focus more on tokens containing fine-grained knowledge, enabling better coarse-grained alignment. Thus we define the importance score as:
(11) $\omega_{t_{i}}=\dfrac{\text{exp}(\text{cos}(z_{t_{i}},z_{t_{0}}))}{\sum_{j=1}^{N}\text{exp}(\text{cos}(z_{t_{j}},z_{t_{0}}))},$
(12) $\omega_{m_{i}}=\dfrac{\text{exp}(\text{cos}(z_{m_{i}},z_{m_{0}}))}{\sum_{j=1}^{N}\text{exp}(\text{cos}(z_{m_{j}},z_{m_{0}}))}.$
where $\omega_{t_{i}}$ is the importance score of word token $t_{i}$, $\omega_{m_{i}}$ is the importance score of motif $m_{i}$. A higher importance score indicates a greater contribution to the global token.
We begin by only pre-training the contrastive alignment task for a few warm-up epochs to ensure valid importance scores. To demonstrate the empirical validity of the importance scores, we summarize the 10 most important and least important tokens in each modality in Tables 1 and Table 2. We observe that the top 10 word tokens are closely associated with specific motifs, indicating their relevance to fine-grained knowledge. In contrast, the bottom 10 word tokens have few associations. Similarly, the top 10 motif tokens often exhibit distinct chemical properties, making them highly meaningful, while the bottom 10 motif tokens are often less meaningful.
Then we jointly pre-train on the contrastive alignment task and masked multimodal modeling task. We mask 15% word tokens and 20% motif tokens based on their importance scores. The probability of masking each token is proportional to its score. As training progresses, the importance scores are dynamically updated, allowing the model to iteratively refine its focus on fine-grained alignment knowledge and improving overall training performance. After pre-training, the final importance scores are shown in Appendix  D.
Table 1. Top 10 and bottom 10 word tokens by average importance score.
Top 10 Bottom 10
Token Score Token Score
triglyceride 0.1419 little 0.002
peptide 0.1290 vegetables 0.002
polysaccharide 0.1284 fruits 0.002
oligo 0.1269 highest 0.002
oligomer 0.1173 game 0.002
anthocyan 0.1053 wild 0.002
diary 0.0998 meat 0.002
isofl 0.0972 chicken 0.002
cannab 0.0953 cheese 0.002
sulfide 0.093 thirty 0.002
Table 2. Top 10 and bottom 10 motif tokens by average importance score.
Top 10 Bottom 10
Token Score Token Score
[NH4+] 0.1766 S 0.0661
[OH-] 0.1588 C 0.0564
C1=CN=CN=C1 0.1285 CCC 0.0521
[NH3+] 0.1252 O 0.0503
Cl 0.123 N 0.0472
O=S 0.119 CCCC(C)OC 0.0407
F 0.112 CCOC(C)O 0.0369
C1=CC=CC=C1 0.108 CCOCC 0.0361
Br 0.1027 C1CCOCC1 0.0299
C1=CNC=C1 0.1016 CCNC(C)=O 0.0231
4. Experiments
In this section, we conduct extensive experiments to demonstrate the effectiveness of FineMolTex. Before evaluating, we first conduct the two pre-training tasks on the PubChemSTM dataset (Liu et al., 2023c), which includes 281K graph-text pairs from PubChem (Kim et al., 2021). Each molecular graph is paired with a textual description that elaborates on its chemical and physical properties or highlights its high-level bioactivities. Details of the pre-training data and process can be found in Appendix C.1.1 and  C.4.
@@ 4. Experiments
4.1. Generalization to Unseen Molecules (RQ1)
Table 3. Accuracy (%±$\pm\sigma$) of graph-text retrieval task on DrugBank-Pharmacodynamics.
@@ 4.1. Generalization to Unseen Molecules (RQ1)
FineMolTex 96.78±$\mathbf{\pm 0.05}$ 92.48±$\mathbf{\pm 0.02}$ 87.94±$\mathbf{\pm 0.14}$ 96.29±$\mathbf{\pm 0.12}$ 91.65±$\mathbf{\pm 0.15}$ 85.07±$\mathbf{\pm 0.11}$
Table 4. Accuracy (%±$\pm\sigma$) of graph-text retrieval task on molecule-ATC.
@@
FineMolTex 76.52±$\mathbf{\pm 0.10}$ 62.75±$\mathbf{\pm 0.06}$ 51.84±$\mathbf{\pm 0.16}$ 76.38±$\mathbf{\pm 0.18}$ 61.72±$\mathbf{\pm 0.09}$ 50.88±$\mathbf{\pm 0.13}$
Table 5. Accuracy (%±$\pm\sigma$) of graph-text retrieval task on DrugBank-Description.
Given Molecular Graph Given Text
T 4 10 20 4 10 20
3D-MoLM 92.81±$\pm 0.23$ 85.71±$\pm 0.19$ 80.20±$\pm 0.33$ 88.31±$\pm 0.32$ 81.23±$\pm 0.07$ 74.40±$\pm 0.39$
MV-Mol 95.13±$\pm 0.16$ 90.28±$\pm 0.21$ 84.83±$\pm 0.34$ 93.54±$\pm 0.32$ 86.58±$\pm 0.52$ 80.75±$\pm 0.43$
FineMolTex 99.60±$\mathbf{\pm 0.06}$ 97.96±$\mathbf{\pm 0.04}$ 96.70±$\mathbf{\pm 0.14}$ 99.62±$\mathbf{\pm 0.02}$ 97.96±$\mathbf{\pm 0.09}$ 96.34±$\mathbf{\pm 0.07}$
@@
Datasets and Baselines. We utilize DrugBank-Pharmacodynamics, molecule-ATC, and DrugBank-Description (Liu et al., 2023c) extracted from the DrugBank database (Wishart et al., 2018) for evaluation. These datasets include molecular graphs and their chemical descriptions. Details of the datasets can be found in Appendix  C.1.2. We compare with five multimodal molecular models: KV-PLM (Zeng et al., 2022), MolCA (Liu et al., 2023a), MoMu-S (Su et al., 2022), MoMu-K (Su et al., 2022), MoleculeSTM (Liu et al., 2023c), 3D-MoLM (Li et al., 2024), and MV-Mol (Luo et al., 2024). Specifically, KV-PLM uses SMILES to represent the structure of the molecule, while others use graph structures.
@@
Results. We report the results on the first two datasets in Tables LABEL:tab:pha,  LABEL:tab:atc, and  LABEL:tab:des. We make the following observations. 1) Across different values of $T$, FineMolTex consistently outperforms the baselines that neglect motif-level knowledge. The superior performance demonstrates that fine-grained motif-level knowledge facilitates generalization to unseen molecules, which likely contain seen motifs. 2) FineMolTex maintains strong performance in both directions (given graph, and given text). The symmetry further indicates that the embedding spaces of both modalities are well-aligned and similarly well-learned. 3) We observe that KV-PLM, which utilizes SMILES to capture molecular structures, is less effective than other models employing graphs, consistent with previous findings (Liu et al., 2023c) that 2D graph structure is more expressive than 1D SMILES.
4.2. Application to Motif-Centered Tasks (RQ2)
Figure 3. Hit ratios of 12 text-based molecule editing tasks.
@@ 4.2. Application to Motif-Centered Tasks (RQ2)
Figure 4. Visual analysis of the output molecules of MoleculeSTM and Motif-MolTex on 4 text-based molecule editing tasks. Differences between the input and output molecule of FineMolTex are highlighted in red and green circles. Lower LogP indicates higher water solubility.
@@
To answer RQ2, we employ a zero-shot text-based molecule editing task, which is highly relevant to practical applications including catalyst design and targeted drug discovery. Specifically, we utilize FineMolTex to collaborate with a molecule generation module, following the design in (Liu et al., 2023c), to modify a specified molecule according to a text prompt. Hence, motif-level knowledge is essential for this task, as the model needs to replace certain motifs with others that are related to specific properties and names as indicated in the text prompt. We defer the technical details to Appendix B. We randomly sample 200 molecules from ZINC (Irwin et al., 2012), and select 12 text prompts, including 8 prompts pertaining to physical properties (Liu et al., 2023c), and 4 based on the names of the motifs. We utilize MoleculeSTM, MoMu, and MolCA as baselines.
@@
Evaluation. We employ different methods to assess whether the generated molecules satisfy the two types of prompts. For the 8 prompts on physical properties, we employ three measures: LogP, QED, and tPSA, which measures solubility (Leo et al., 1971), drug-likeness (Bickerton et al., 2012), and permeability (Ertl et al., 2000), respectively. We consider the editing to be successful if the difference in measurements between the input and output molecules exceeds a specified threshold $\Delta$, which we have set to 00 following one of the settings in literature (Liu et al., 2023c). For the 4 prompts based on motif names, we use RDKit (Landrum, 2024) to verify the presence of the indicated motifs. For all 12 prompts, we report the hit ratio: the proportion of generated molecules that meet our expectations.
@@
Results. As shown in Figure 3, FineMolTex shows superior performance on these prompts, especially on the 4 prompts with motif names. Notably, we achieve a relative gain of up to 230% over the best-performing baseline, demonstrating that FineMolTex has an advanced understanding of motif-level knowledge. We also visualize the output molecules of MoleculeSTM and FineMolTex in Figure 4. It can be observed that while MoleculeSTM produces incorrect molecules, FineMolTex accurately generates the intended molecules. Specifically, when prompted to generate molecules that are soluble in water, FineMolTex successfully creates molecules with lower LogP than the input molecule, as lower LogP indicates higher water solubility. Similarly, when prompted to generate molecules with hydroxyl groups or chlorine atoms, FineMolTex correctly does so. These results confirm that FineMolTex possesses a deeper understanding of motif-level knowledge, thereby enhancing the generative capabilities.
4.3. Application to Single-Modality Task (RQ3)
@@
Table 6. Downstream results (%±$\pm\sigma$) on eight binary classification datasets from MoleculeNet.
@@
MolCA 70.0$\pm$0.5 77.2±$\mathbf{\pm 0.5}$ 64.5$\pm$0.8 63.0$\pm$1.7 89.5$\pm$0.7 72.1±$\pm 1.3$ 77.2±$\pm 0.6$ 79.8$\pm$0.5 74.2
MoleculeSTM 70.0±$\pm 0.5$ 76.9±$\pm 0.5$ 65.1±$\pm 0.4$ 61.0±$\pm 1.1$ 92.5±$\pm 1.1$ 73.4±$\pm 2.9$ 77.0±$\pm 1.8$ 80.8±$\pm 1.3$ 74.6
FineMolTex 73.5±$\mathbf{\pm 1.6}$ 77.1±$\pm 1.2$ 68.6±$\mathbf{\pm 0.9}$ 64.8±$\mathbf{\pm 1.4}$ 92.5±$\mathbf{\pm 0.8}$ 76.3±$\mathbf{\pm 1.2}$ 79.0±$\mathbf{\pm 1.4}$ 84.0±$\mathbf{\pm 1.5}$ 76.9
Table 7. Ablation study (%±$\pm\sigma$) on molecule-ATC and DrugBank-Pharmacodynamics.
@@
w/o mmm 68.85±$\pm 0.32$ 69.34±$\pm 0.14$ 90.18±$\pm 0.08$ 90.52±$\pm 0.14$
motif mask only 72.64±$\pm 0.05$ 70.96±$\pm 0.20$ 92.24±$\pm 0.12$ 92.06±$\pm 0.31$
word mask only 73.68±$\pm 0.11$ 72.47±$\pm 0.09$ 93.28±$\pm 0.26$ 92.97±$\pm 0.17$
w/o cross-attention 69.92±$\pm 0.25$ 69.35±$\pm 0.22$ 92.66±$\pm 0.08$ 92.85±$\pm 0.24$
random mask only 73.95±$\pm 0.15$ 73.34±$\pm 0.19$ 93.57±$\pm 0.32$ 93.25±$\pm 0.18$
FineMolTex 76.52±$\mathbf{\pm 0.10}$ 76.38±$\mathbf{\pm 0.18}$ 96.78±$\mathbf{\pm 0.05}$ 96.29±$\mathbf{\pm 0.12}$
While FineMolTex can simultaneously utilize pre-trained knowledge from both graphs and texts, we also verify its effectiveness on single-modality tasks, namely, molecular property prediction tasks. We use MoleculeNet (Wu et al., 2018) as the dataset, which only provides molecular graphs as input without texts. More specifically, there are eight binary classification tasks, and we report ROC-AUC for evaluation. More detailed dataset descriptions are provided in Appendix C.1.3.
Baselines. We compare FineMolTex against nine baselines, including 1) five pre-trained GNN models: AttrMasking (Hu et al., 2019), ContextPred (Hu et al., 2019), InfoGraph (Sun et al., 2019), MolCLR (Wang et al., 2021b), and GraphMVP (Liu et al., 2019b); 2) three graph-text multimodal models: MoMu-S (Su et al., 2022), MoMu-K (Su et al., 2022), and MoleculeSTM (Liu et al., 2023c), and 3) one SMILES-text multimodal model: KV-PLM (Zeng et al., 2022).
Results. As shown in Table  6, FineMolTex consistently outperforms all baselines, achieving relative gains of 3.2%, 2.4%, and 4.7% on SIDER, MUV, and BACE, respectively, compared to the best baseline. The promising performance of FineMolTex indicates that it implicitly utilizes pre-trained knowledge from the text modality even when the input consists solely of graphs. Additionally, KV-PLM exhibits a notable performance gap from other models, due to its use of 1D SMILES strings for molecular structure and a smaller pre-training dataset.
4.4. Analysis of Learned Fine-grained Knowledge (RQ4)
Visualization of Motif and Word Embeddings. To evaluate whether FineMolTex captures fine-grained alignment knowledge, we select motif and word tokens from 1,500 graph-text pairs in the PubChemSTM dataset, excluding meaningless words such as “this” and “a”. In total, we visualize 3,469 motif tokens and 6,914 text tokens with $t$-SNE (Maaten and Hinton, 2008) in Figure  5, where triangles denote text tokens, and circles denote motif tokens, with different colors indicating various labels. To examine the details of the tokens, we zoom into several regions in the figure, retaining only the colors and legends of the tokens we are interested in. For brevity, we utilize SMILES to represent the motif structures. We observe that text and motif tokens corresponding to each other are also close in the embedding space. For instance, in the pink frame, the word “ammonium” is close to the motif tokens “[NH2+]=O”, “C=[NH2+]”, and “[NH3+]O”, which are related to “ammonium.” In the blue frame, the word “poison” is adjacent to the motifs “[AsH3]” and “O[AsH2]”, which are poisonous. In the orange frame, the word “sulf” is close to the motif tokens “OS”, “CCSSCC”, and “CC1=CSC(C)=N1”, all of which represent sulfides. These results demonstrate that FineMolTex learns the connections between motifs and their chemical names and properties, thereby significantly enhancing its expressiveness.
Figure 5. Visualization of motif tokens and word tokens using $t$-SNE. Triangles denote word tokens; circles denote motif tokens.
Figure 6. Explaination of the prediction of certain masked motifs based on text tokens utilizing LIME.
Predictions Based on Fine-grained knowledge. To further verify that FineMolTex can utilize the learned fine-grained alignment knowledge for predictions, we utilize Local Interpretable Model-Agnostic Explanation (LIME) (Ribeiro et al., 2016), a well-established tool that can explain the predictions of certain masked motifs based on text tokens. By perturbing the input text, LIME observes how the model’s predictions change with variations in the input text. Then, LIME fits these perturbed texts and the prediction results to an interpretable model such as a linear model. This approach allows us to quantify the significance of each text token in predicting the motifs, thereby revealing the fine-grained knowledge learned by FineMolTex. The results are shown in Figure 6, where text tokens with higher interpretive weights are more crucial for predictions, and thus more relevant to the masked motifs. Specifically, the word with the highest interpretive weights in (1) is “heteroarene”, which directly refers to the name of the masked motif. In (2), the word “phenyl” refers to the masked motif. In (3), the masked motif makes the molecule “bromide”. These findings demonstrate that FineMolTex has effectively acquired motif-level knowledge.
Additional analysis such as fine-grained alignment information in the joint embedding space can be found in Appendix  D.3
4.5. Ablation Study for Masking and Cross-Attention Layers (RQ5)
To thoroughly explore the impact of the key components in FineMolTex, we compare to several variants, including w/o mmm, which drops the masked multimodal modeling task altogether; motif mask only, which only mask motif tokens; word mask only, which only mask word tokens; w/o cross-attention, which excludes cross-attention layers. random mask only, which masks motif or word tokens randomly rather than based on importance. We evaluate these variants on the graph-text retrieval task used in RQ1, on two datasets with $T=4$. As reported in Table LABEL:tab:abla, FineMolTex consistently surpasses the other variants. Specifically, without the masked multimodal modeling task, “w/o mmm” fails to capture fine-grained knowledge at all, resulting in the poorest performance. “motif mask only” and “word mask only” outperform “w/o mmm,” because they still enable some level of fine-grained knowledge learning by either predicting motifs based on word tokens or predicting word tokens based on motif tokens. However, they are less effective than FineMolTex, which masks both words and motifs for mutual alignment. Lastly, without the cross-attention layers, “w/o cross-attention” cannot integrate token embeddings from different modalities, without importance-based masking, “random mask only” cannot effectively focus on the tokens that contain fine-grained knowledge, thereby hampering their ability to effectively learn fine-grained knowledge. These observations demonstrate the effectiveness of each component.
Additional experimental results, including further ablation studies and a comparison of pre-training and inference times, are provided in Appendix  D.
5. Conclusions